# Susceptibility for extremely low external fluctuations and critical behaviour of Greenberg-Hastings neuronal model

### v1

@@ V Discussion
The usage of three different calculation methods allowed us to obtain reliable estimations of the critical exponents, which do not seem to agree (to the best of our knowledge) with any known universality class. The last fact is probably associated with the presence of quenched disorder in the model, since rare regions effects [8, 19] in this case cannot be discarded; future works along this line will include critical exponents calculations varying the disorder, either in the synapsis probability distribution or in the WS rewiring probability. Another extension of the present work could be to perform a finite size scaling analysis in the $r_{1}\to 0$ limit of the variability in the lasting time activity when starting from a single neuron activation, as done in Ref.[1] for a fixed system size (See Fig.1 in the Supplemental Material).
@@
We also analysed the influence of different activation mechanisms around the activation transition point by comparing numerical simulations with a mean field approximation that neglects cooperative activations. We found that cooperative mechanisms become progressively relevant as the average degree of the network increases, being almost negligible for small values of $\langle k\rangle$ and dominant for large values of it. In particular, in the range of values of $\langle k\rangle$ where the activation transition is continuous, those mechanisms contribute to the phenomenon only for intermediate values of the threshold $T$. Consequently, the mean field approximation works well for low and high values of $T$, including the critical region. In particular, the prediction of a logarithmic increase of the critical value $T_{c}$ with $\langle k\rangle$ when $r_{1}=0$, agrees very well with the numerical results in almost the whole region of continuous phase transitions $4\leq\langle k\rangle\leq 24$. Our mean field approximation also predicts a critical phenomenon in the directed percolation universality class, in agreement with the prediction of a related model without cooperative interactions at all [7]. Also, the result $f_{a}(T=T_{c},r_{1})\sim r_{1}^{1/\delta_{MF}}$ within the mean field approximation reinforces the idea that $r_{1}$ acts as an external field conjugated to $f_{a}$. Finally, in the region $\langle k\rangle>24$, where the dynamical phase transition becomes discontinuous and is dominated by the cooperative activation mechanism, $T_{c}$ grows linearly with $\langle k\rangle$, as expected.

### v2

@@ V Discussion
The usage of three different calculation methods allowed us to obtain reliable estimations of the critical exponents, which do not seem to agree (to the best of our knowledge) with any known universality class. The last fact is probably associated with the presence of quenched disorder in the model, since rare regions effects [8, 20] in this case cannot be discarded (a possibility also supported by the apparent presence of power law behaviours for many different values of $T$ shown in the supplementary material); future works along this line will include critical exponents calculations varying the disorder, either in the synapsis probability distribution or in the WS rewiring probability. Another extension of the present work could be to perform a finite size scaling analysis in the $r_{1}\to 0$ limit of the variability in the lasting time activity when starting from a single neuron activation, as done in Ref.[1] for a fixed system size.
@@
We also analysed the influence of different activation mechanisms around the activation transition point by comparing numerical simulations with a mean field approximation that neglects cooperative activations. We found that cooperative mechanisms become progressively relevant as the average degree of the network increases, being almost negligible for small values of $\langle k\rangle$ and dominant for large values of it. In particular, in the range of values of $\langle k\rangle$ where the activation transition is continuous, those mechanisms contribute to the phenomenon only for intermediate values of the threshold $T$. Consequently, the mean field approximation works well for low and high values of $T$, including the critical region. In particular, the prediction of a logarithmic increase of the critical value $T_{c}$ with $\langle k\rangle$ when $r_{1}=0$, agrees very well with the numerical results in almost the whole region of continuous phase transitions $4\leq\langle k\rangle\leq 24$. These results are consistent with those of a recent work[22], where an equation similar to Eq.(8) is derived under the same assumption that single activations are dominant. The validity of this assumption was supported by both theoretical and computational evidence for a wide range of threshold values in different network topologies and weight distributions. Moreover, they also show that cooperative activation becomes more relevant as $\langle k\rangle$ increases, consistently with our results.
Our mean field approximation also predicts a critical phenomenon in the DP universality class, in agreement with the prediction of a related model without cooperative interactions at all [7]. However, the critical exponents for the model defined on a WS network obtained numerically do not agree with any known universality class, particularly with the mean field DP ones that would be expected in this case. This result could be consistent with the fact that the hyperscaling relation seems not to be satisfied. However, the numerical evidence of rapidity reversal symmetry in the model (see supplementary material) makes the whole scenario inconclusive, leaving the question about universality in the model defined on a WS network open.
The result $f_{a}(T=T_{c},r_{1})\sim r_{1}^{1/\delta_{h,MF}}$ within the mean field approximation reinforces the idea that $r_{1}$ acts as an external field conjugated to $f_{a}$. While the numerical results in a WS network also support this conclusion, they are consistent with a threshold dependent critical exponent $\delta_{h}(T)$ for a wide range of values $T\geq T_{c}$, thus providing another evidence of the presence of rare-regions effects.
Finally, in the region $\langle k\rangle>24$, where the dynamical phase transition becomes discontinuous and is dominated by the cooperative activation mechanism, $T_{c}$ grows linearly with $\langle k\rangle$, as expected.